OpenStudy (anonymous):

Let L= the following vectors and let V = span{L}. Find a basis of V contained in L.

OpenStudy (anonymous):

\[\left\{ \left(\begin{matrix}1 \\ 1\\-1\end{matrix}\right),\left(\begin{matrix}1 \\ 1\\11\end{matrix}\right),\left(\begin{matrix}1 \\ 1\\2\end{matrix}\right),\left(\begin{matrix}2 \\ 4\\1\end{matrix}\right) \right\}\]